OpenStudy (anonymous):

can someone please help me find the range of this function f(x) = (x - 2)2 + 2

OpenStudy (gorv):

range means what output we will get from this

TheSmartOne (thesmartone):

If you look at its grapg you see the range starts from 2 and goes to positve infinity...

OpenStudy (gorv):

whatever value of x we will put here that will give positive output

OpenStudy (gorv):

beacuse square make that positive

OpenStudy (gorv):

min value we will get at x=2

TheSmartOne (thesmartone):

So in interval notion that would be \[[2,\infty)\]

OpenStudy (anonymous):

thank you i think i got it now

TheSmartOne (thesmartone):

np

OpenStudy (anonymous):

if f(x) = (x-2)2+2 then f(x) = (x^2) - 4x +6 With y intercept being 6 Range: All real numbers greater than 6 since the parabola's minimum value is 6. If you make a graph, you can clearly see it.